OpenStudy (anonymous):

Which of the following shows the square completed in this equation? x2 + 3x - 5 = 0 I dont get how to do this one.

OpenStudy (anonymous):

which of the following?

OpenStudy (anonymous):

it should look something like \[(x+\frac{3}{2})^2=\text{something}\]

OpenStudy (anonymous):

\[x^2+3x-5=0\] \[x^2+3x=5\] \[(x+\frac{3}{2})^2=5+\frac{9}{4}=\frac{29}{4}\] \[x+\frac{3}{2}=\pm\frac{\sqrt{29}}{2}\] \[x=-\frac{3}{2}\pm\frac{\sqrt{29}}{2}\]

OpenStudy (anonymous):

thanks! but which one wuld it be? (x + 9)2 = 14 (x - 3)2 = 8 (x + 1.5)2 = 7.25

OpenStudy (mertsj):

Did you see that Satellite said the left side should be (x+3/2)^2

OpenStudy (anonymous):

Yes. so it would be the 2nd one? sorry just confused

OpenStudy (anonymous):

It's the third one. You get there after this stage: (x+3/2)^2=5+9/4 3/2 is 1.5 and 5+9/4 is 7.5

OpenStudy (anonymous):

thanks!